I really need your help in solving this problem. I appreciate your time and efforts. Have a terrific day :) Here is the three part question?
One sample has a mean of M= 4 and second sample has M=8. The two samples are combined into a single set of scores.
A. What is the mean for the combined set if both of the Original Samples have n = 7 scores?
B. What is the mean for the combined set if the first sample has n = 3 and the second sample has n = 7?
C. What is the mean for the combined set if the first sample has n =7 and the second sample has n = 3?

Answer by macston(5194)

Part A: First sample with M=4 and 7 scores Second sample M=8 with 7 scores
With average (mean)=4 with 7 scores would sum (4 x 7)=28
With average (mean)=8 with 7 scores would sum (8 x 7)=56
total points=28+56=84 total scores (7+7)=14 scores
so combined mean is 84 total points/14 scores=combined M=6 ANSWER A Mean=6
Part B:
With average (mean)=4 with 3 scores would sum (4 x 3)=12
With average (mean)=8 with 7 scores would sum (8 x 7)=56
Total points=12+56=68; total scores (3+7)=10 scores
so combined mean is 68 total points/10 scores=combined M=6.8 ANSWER B Mean=6.8
Part C:
With average (mean)=4 with 7 scores would sum (4 x 7)=28
With average (mean)=8 with 3 scores would sum (8 x 3)=24
Total points=28+24=52; total scores (7+3)=10 scores
so combined mean is 52 total points/10 scores=combined M=5.2 ANSWER C Mean=5.2